Find all real values of k for which 2x squared plus kx plus 3 equals 0 has a double root.
Answer by stanbon(75887)

Find all real values of k for which 2x squared plus kx plus 3 equals 0 has a double root.
-----
Note: Every quadratic has 2 roots. By a "double root"
I assume you meed a root with multiplicity 2.
----------------
2x^2 + kx + 3 = 0
A double root means b^2-4ac = 0
------
k^2-4*2*3 = 0
k^2 = 24
k = 2sqrt(6)
===================
Cheers,
Stan H.
===================